if b is 1.2 and t is 4. first make x the subject and work out its value in this equation
b = log4x - 2log t

Answer by stanbon(75887)

if b is 1.2 and t is 4. first make x the subject
and work out its value in this equation
------------------------
b = log4x - 2log t
---
1.2 = log4(x) - 2log(4)
---
1.2 = log4(x) - log(16)
----
1.2 = log4(x) - 1.2041
----
log4(x) = 2.4041
---
x = 4^2.4041
---
x = 28.016..
===============
Cheers,
Stan H.
============